A=9 C=5 B=120 degrees find the measure of angle A
Answer by stanbon(75887)

A=9 C=5 B=120 degrees find the measure of angle A
--------
Draw a 120 degree angle; Let the sides be A = 9 and C = 5
Find side B using Law of Cosines::
B^2 = A^2 + C^2- 2A*C*cos(120)
B^2 = 81 + 25 - 2*9*5*(-1/2)
-----
B^2 = 151
B = 12.29
---------------------
Find angle A using Law of Cosines::
cos(A) = [B^2+C^2-A^2]/(2B*C)
cos(A) = [151+25-81]/(2*12.29*5) = 0.7730
-------------
angle A = arcCos(0.7730) = 39.38 degrees
-------------
Cheers,
Stan H.